OpenStudy (anonymous):

Solve the trigonometric equation 3tan^2 2theta - 1 = 0

OpenStudy (anonymous):

Interval [0, pi]

OpenStudy (anonymous):

theta=x 3tan^2 (2x) = 1 tan^2(2x) = 1/3 tan(2x) = 1/rt(3)........rt=root under tan(2x) = tan 30 2x = 30 x= 15